. Calculate the percent by mass of sodium hypochorite in her disinfectant solution. (


5Tbsp. of NaClO (sodium hypochlorite) and 3.8 liters of H2O (water). In units of grams


for both of the solute and solvent.

1
Expert's answer
2022-04-29T01:08:06-0400

5 Tbsp of NaClO is nearly 41.4 g.

Density of water is nearly 1 g/ml, so 3.8 l of water is 3800 g.

"w(NaClO)=\frac{41.4x100}{41.4+3800}=1.08\%"

"w(H_2O)=\frac{3800x100}{41.4+3800}=98.92\%"